If a perfectly competitive firm is producing 150 units of output at a price of P=$20, where the MC of the 150th unit of output is MC=$20, the ATC of the 150th unit is ATC=$10, and the AVC of the 150th unit is AVC=$8, then which of the following statements is not correct?

a. The firm should shut down when the price is less or equal to $8.

b. The firm is producing at the profit maximizing level of output.

c. The firm is earning a profit of $2000.

d. More firms will want to enter the market.
Answer #1

Ans) the correct option is c. The firm is earning a profit of $2000.

At Profit maximization, P = MC

So profit = ( P - ATC) * Q = (20 - 10) * 150 = 1500. Since the Profits are positive, firms should enter the market